OpenStudy (anonymous):

The morning temperature in Dallas is 47 degrees and rises 10 degrees per hour. The morning temperature in Miami is 67 degrees and rises 5 degrees per hour. In how many hours will the two cities be the same temperature

OpenStudy (johnweldon1993):

So you set up 2 equations and set them equal to each other 47 + 10h = 67 +5h 5h = 20 h = 4 in 4 hours Dallas will be 87 and Miami will be 87 as well

OpenStudy (anonymous):

thank you!

OpenStudy (johnweldon1993):

No problem :)